# Topology-Preserving Line Densification for Creating Contiguous Cartograms

### v1

@@ 1 Introduction
(a) Equal-Area Map
(b) Cartogram
Figure 1: Maps of Belgium. (a) Belgium’s regions on an equal-area map, abbreviated as BRU (Brussels), VLG (Flanders), and WAL (Wallonia). (b) Contiguous cartogram representing population. As Brussels is the most densely populated region, its area is enlarged in the cartogram.
@@
To evaluate the proposed densification algorithm’s effectiveness, we augment the flow-based cartogram generation algorithm by Gastner et al. (2018) with our densification method. We refer to the resulting method as 5FCarto, where “5F” refers to five key properties: flow-based, faultless (i.e., no topology violations), free (open-source under GPL v3), fast (quicker than other current methods), and focused (primarily allocating computational resources to regions that require significant transformation).
@@
Figure 2: Illustration of line densification when approximating a curve with a finite number of points. Panel (a) depicts the ground truth, showing the precise geometry of two curves, $C_{1}$ and $C_{2}$. Panel (b) illustrates finite approximations, $A_{1}$ and $A_{2}$, that use a sparse distribution of points along the respective curves. Topology is violated as $A_{1}$ intersects itself and also $A_{2}$. Panel (c) demonstrates line densification resulting in $A_{1}^{\prime}$ and $A_{2}^{\prime}$ preserving the true topology.
@@ 3. P3:
Figure 3: Illustration of a topology violation caused by projecting polylines. The figure depicts two polylines, one colored blue (ABC) and the other red (DE). Panel (a) illustrates the original geometry, while panel (b) shows the situation after transformation via a density-equalizing map projection. The transformed black grid lines indicate the projection. The projected positions of the polyline vertices A to E are interpolated on the basis of the quadrilaterals formed by the projected grid lines in panel (b). Although the blue and red polylines do not intersect before the transformation, they cross each other afterward, resulting in an invalid topology.
@@ 4.1 Quadtree Construction
Figure 4: Illustration of the quadtree splitting criteria. The squares that are delineated by thin gray lines indicate a rasterized density grid, while those outlined in black denote quadtree leaf nodes. The quadtree is constructed by recursively splitting an $n\times n$ cell containing the largest difference between maximum and minimum density into four $(n/2)\times(n/2)$ cells until a predetermined number of leaf nodes (7 in this example) is reached. The squares in the raster grid that contain the maximum and minimum density values are annotated with those values. Panel (a) depicts the initial $4\times 4$ cell, which is split due to having the largest internal density difference (22) and because the current number of leaf nodes (1) has not reached the target (7). Panel (b) illustrates the resulting subdivision with 4 leaf nodes, where again the node with the largest density difference (12, bottom-right) is selected for further splitting, as the target number of leaf nodes has still not been met. Panel (c) shows the final subdivision with 7 leaf nodes, at which point the desired number of leaf nodes has been reached, marking the completion of the quadtree.
@@ 4.3 Line Densification
After the constrained Delaunay triangulation $T$ is created, we insert points at all intersections of polygon edges with any Delaunay triangle boundary (Figure 6(a)). The projected coordinates of the original and added points are then interpolated using the uniquely defined piecewise affine transformation that maps the unprojected Delaunay triangles to their respective projections, as shown in Figure 6(b).
@@
Figure 5: Illustration of the triangulation process. (a) Original graded quadtree, as obtained in Figure 4(c). (b) Quadtree corners projected according to the density grid from Figure 4. The projected quadtree cell edges (black solid lines) serve as constraints for computing a Delaunay triangulation, $T_{\text{temp}}$, of the projected corners. The additional triangulation edges are represented by dashed lines. (c) After unprojecting the edges, the final constrained Delaunay triangulation, $T$, comprises both the solid black lines, as well as the red, green and blue dashed lines. Although the yellow dashed line represents the Delaunay triangulation of the transformed quadtree, we do not insert it into the original quadtree because it overlaps with a boundary of a quadtree leaf node. Thus, we initially only add the blue dashed line as a constraint in the upper right leaf node. Subsequently, the green dashed line is added to complete the Delaunay triangulation under the specified constraints.
@@ 4.4 Guaranteed Topology Preservation
To establish that 5FCarto guarantees topologically valid projected polygons, we first recall that density-equalizing map projections are invertible, as explained in Section 2. Assuming that the Jacobian is sufficiently smooth to preserve orientation in the projected triangulation, the affine mappings on adjacent Delaunay triangles align continuously, preventing any intersections along their shared edges. Furthermore, line segments inside the same triangle will not intersect after the projection because the affine transformation of the triangle has a nonzero determinant, establishing properties P1 and P2. Finally, because property P3 follows directly from the orientation‑preserving nature of the triangulation, the projected polygons are topologically valid.
Figure 6: Illustration of our algorithm designed to prevent polyline intersections. (a) The positions of the polyline vertices are identical to those in Figure 3(a); the quadtree and the Delaunay triangulation are those shown in Figure 5(c). Additionally, polyline vertices, represented by open circles, are added at intersections with the Delaunay triangle boundaries. (b) When the vertices added in (a) are transformed using the same density-equalizing map projection as in Figure 3(b), alongside vertices A to E, the red and blue polylines do not intersect. Thus, in contrast to the situation depicted in Figure 3(b), the original topology is preserved.
Figure 7: Illustration of Belgium’s constrained Delaunay triangulation. Among the various possible triangulations of the quadtree corners in panel (a), the selected triangulation maximizes the minimum angle of the projected triangles in panel (b).
@@ 5 Evaluation
(a) Equal-Area Map
(b) F4Carto Cartogram (Error: 29.719)
(c) 5FCarto Cartogram (Error: 0.007)
(d) BFB Cartogram (Error: 2.388)
Figure 8: Illustration of Singapore’s equal-area geometry, divided into planning areas, and population cartograms generated by F4Carto, 5FCarto, and BFB. The values in parentheses indicate the maximum relative area error for each cartogram. (a) Equal-area map of Singapore, provided as input. (b) Cartogram generated by F4Carto, with 483 overlapping intersections indicated by blue semicircles. This cartogram exhibits a high maximum relative area error, and it produces nearly rectilinear boundaries that deviate significantly from the true geographic contours. (c) Cartogram generated by 5FCarto, with no intersections present. (d) Cartogram generated by BFB, displaying 316 self-intersections (red upper semicircles) and 423 overlapping intersections (blue lower semicircles).
@@ 5.1.3 Deformation
To extend these metrics to an entire map or cartogram, which is composed of multiple regions, we compute each metric for every corresponding pair of regions between the original map and the cartogram. Each region may consist of multiple disjoint polygons with holes. When there is a one‑to‑one relationship between the polygon rings of the equal‑area map (both outer boundaries and holes) and those of the cartogram, deformation metrics can be computed for each polygon. However, since F4Carto sometimes removes small polygons and polygons with holes, a one-to-one matching between the disjoint polygons becomes infeasible. Consequently, for the Fréchet and Hausdorff distances, we only consider the polygon with holes having the largest area for each region. We then use the outer ring from both the equal-area map and the cartogram for calculating these distances. In contrast, no such adjustment is necessary for the symmetric difference, allowing the use of all available polygons with holes.
@@ 6 Application: 2024 Presidential Election Cartogram of the United States
As a real-world case study, we apply the 5FCarto algorithm proposed in this paper to data from the 2024 United States presidential election. These election results have formed the basis for cartograms published by various news outlets—for example, in articles by ABC News (2024), Levinson et al. (2024), and Breteau and Ferrer (2024)—because such maps eliminate visual overemphasis of land area relative to electoral weight. Figure 9 shows the equal-area reference map used as input to 5FCarto in panel (a) and the resulting cartogram in panel (b), where areas of all states are proportional to their number of electoral votes.
@@
Figure 9: Maps of the United States. Colors represent the party that won the electoral votes in the 2024 presidential election—by state or, in Maine (ME) and Nebraska (NE), congressional district. (a) Equal-area map. (b) Cartogram generated by 5FCarto, with each region’s area proportional to its electoral-vote count.
@@ Acknowledgements
This research is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Award Number: MOE-T2EP20222-0039).
@@
No potential conflict of interest was reported by the author(s).
@@
The data and code that support the findings of this study are openly available at https://github.com/mgastner/cartogram_generator and in the supplementary information.

### v2

@@ 1 Introduction
Figure 1: Maps of Belgium. a) Belgium’s regions on an equal-area map, abbreviated as BRU (Brussels), VLG (Flanders), and WAL (Wallonia). b) Contiguous cartogram representing population. As Brussels is the most densely populated region, its area is enlarged in the cartogram.
@@
To evaluate the proposed densification algorithm’s effectiveness, we augment the flow-based cartogram generation algorithm by Gastner et al. (2018) with our densification method. We refer to the resulting method as 5FCarto, where “5F” refers to five key properties: flow-based, faultless (i.e., no topology violations), free (open-source under AGPL-3.0), fast (quicker than other current methods), and focused (primarily allocating computational resources to regions that require significant transformation).
@@
Figure 2: Illustration of line densification when approximating a curve with a finite number of points. Panel a) depicts the ground truth, showing the precise geometry of two curves, $C_{1}$ and $C_{2}$. Panel b) illustrates finite approximations, $A_{1}$ and $A_{2}$, that use a sparse distribution of points along the respective curves. Topology is violated as $A_{1}$ intersects itself and also $A_{2}$. Panel c) demonstrates line densification resulting in $A_{1}^{\prime}$ and $A_{2}^{\prime}$ preserving the true topology.
@@ 3. P3:
Figure 3: Illustration of a topology violation caused by projecting polylines. The figure depicts two polylines, one colored blue (ABC) and the other red (DE). Panel a) illustrates the original geometry, while panel b) shows the situation after transformation via a density-equalizing map projection. The transformed black grid lines indicate the projection. The projected positions of the polyline vertices A to E are interpolated on the basis of the quadrilaterals formed by the projected grid lines in panel b). Although the blue and red polylines do not intersect before the transformation, they cross each other afterward, resulting in an invalid topology.
@@ 4.1 Quadtree Construction
Figure 4: Illustration of the quadtree splitting criteria. The squares that are delineated by thin gray lines indicate a rasterized density grid, while those outlined in black denote quadtree leaf nodes. The quadtree is constructed by recursively splitting an $n\times n$ cell containing the largest difference between maximum and minimum density into four $(n/2)\times(n/2)$ cells until a predetermined number of leaf nodes (7 in this example) is reached. The squares in the raster grid that contain the maximum and minimum density values are annotated with those values. Panel a) depicts the initial $4\times 4$ cell, which is split due to having the largest internal density difference (22) and because the current number of leaf nodes (1) has not reached the target (7). Panel b) illustrates the resulting subdivision with 4 leaf nodes, where again the node with the largest density difference (12, bottom-right) is selected for further splitting, as the target number of leaf nodes has still not been met. Panel c) shows the final subdivision with 7 leaf nodes, at which point the desired number of leaf nodes has been reached, marking the completion of the quadtree.
@@ 4.3 Line Densification
After the constrained Delaunay triangulation $T$ is created, we insert points at all intersections of polygon edges with any Delaunay triangle boundary (Figure 6a). The projected coordinates of the original and added points are then interpolated using the uniquely defined piecewise affine transformation that maps the unprojected Delaunay triangles to their respective projections, as shown in Figure 6b.
@@
Figure 5: Illustration of the triangulation process. a) Original graded quadtree, as obtained in Figure 4. b) Quadtree corners projected according to the density grid from Figure 4. The projected quadtree cell edges (black solid lines) serve as constraints for computing a Delaunay triangulation, $T_{\text{temp}}$, of the projected corners, with additional triangulation edges represented by dashed lines. c) After unprojecting the edges, the final constrained Delaunay triangulation, $T$, comprises both the solid black lines and the red, green, and blue dashed lines. Although the yellow dashed line represents the Delaunay triangulation of the transformed quadtree, we do not insert it into the original quadtree because it overlaps with a boundary of a quadtree leaf node. Thus, we initially only add the blue dashed line as a constraint in the upper right leaf node before adding the green dashed line to complete the triangulation under the specified constraints.
@@ 4.4 Guaranteed Topology Preservation
To establish that 5FCarto guarantees topologically valid projected polygons, we first recall that density-equalizing map projections are invertible, as explained in Section 2. Assuming that the Jacobian is sufficiently smooth to preserve orientation in the projected triangulation, the affine mappings on adjacent Delaunay triangles align continuously, preventing any intersections along their shared edges. Furthermore, line segments inside the same triangle will not intersect after the projection because the affine transformation of the triangle has a nonzero determinant, establishing properties P1 and P2. Finally, because property P3 follows directly from the orientation–preserving nature of the triangulation, the projected polygons are topologically valid.
Figure 6: Illustration of our algorithm designed to prevent polyline intersections. a) The positions of the polyline vertices are identical to those in Figure 3; the quadtree and the Delaunay triangulation match those in Figure 5. Additional polyline vertices, represented by open circles, are added at intersections with the Delaunay triangle boundaries. b) When the vertices added in a) are transformed using the same density-equalizing map projection as in Figure 3, alongside vertices A to E, the red and blue polylines do not intersect. Thus, in contrast to the earlier situation, the original topology is preserved.
Figure 7: Illustration of Belgium’s constrained Delaunay triangulation. Among the various possible triangulations of the quadtree corners in panel a), the selected triangulation maximizes the minimum angle of the projected triangles in panel b).
@@ 5 Evaluation
Figure 8: Illustration of Singapore’s equal-area geometry, divided into planning areas, and population cartograms generated by F4Carto, 5FCarto, and BFB. The values in parentheses indicate the maximum relative area error for each cartogram. a) Equal-area map of Singapore, provided as input. b) Cartogram generated by F4Carto, with 483 overlapping intersections indicated by blue semicircles. This cartogram exhibits a high maximum relative area error, and it produces nearly rectilinear boundaries that deviate significantly from the true geographic contours. c) Cartogram generated by 5FCarto, with no intersections present. d) Cartogram generated by BFB, displaying 316 self-intersections (red upper semicircles) and 423 overlapping intersections (blue lower semicircles).
@@ 5.1.3 Deformation
To extend these metrics to an entire map or cartogram, which is composed of multiple regions, we compute each metric for every corresponding pair of regions between the original map and the cartogram. Each region may consist of multiple disjoint polygons with holes. When there is a one–to–one relationship between the polygon rings of the equal-area map (both outer boundaries and holes) and those of the cartogram, deformation metrics can be computed for each polygon. However, since F4Carto sometimes removes small polygons and polygons with holes, a one–to–one matching between the disjoint polygons becomes infeasible. Consequently, for the Fréchet and Hausdorff distances, we only consider the polygon with holes having the largest area for each region. We then use the outer ring from both the equal–area map and the cartogram for calculating these distances. In contrast, no such adjustment is necessary for the symmetric difference, allowing the use of all available polygons with holes.
@@ 6 Application: 2024 Presidential Election Cartogram of the United States
As a real-world case study, we apply the 5FCarto algorithm proposed in this paper to data from the 2024 United States presidential election. These election results have formed the basis for cartograms published by various news outlets—for example, in articles by ABC News (2024), Levinson et al. (2024), and Breteau and Ferrer (2024)—because such maps eliminate visual overemphasis of land area relative to electoral weight. Figure 9 shows the equal-area reference map used as input to 5FCarto in panel a) and the resulting cartogram in panel b), where areas of all states are proportional to their number of electoral votes.
@@
Figure 9: Maps of the United States. Colors represent the party that won the electoral votes in the 2024 presidential election—by state or, in Maine (ME) and Nebraska (NE), congressional district. a) Equal-area map. b) Cartogram generated by 5FCarto, with each region’s area proportional to its electoral-vote count.
@@ Acknowledgements
This project is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (EP2) programme (Award No. MOE-T2EP20221-0007). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of the Ministry.
@@
The authors report that there are no competing interests to declare.
@@
The data that support the findings of this study are openly available in figshare at https://doi.org/10.6084/m9.figshare.28478408.